# The fractional $p$-Laplacian on hyperbolic spaces

### v2

@@ 5. Pointwise convergence
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}G_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &\leq(1+\varepsilon)^{2m+1}(p+2m+1)\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p+2m}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}(p+2m+1)\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &=\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}(p+2m+1)\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\\ &=\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}\frac{2^{m}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+3}{2}\right).\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\sqrt{\frac{\pi}{2}}\frac{2^{m-1}}{p}\left(\frac{2}{m-1/2}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m}{2}\right).\end{split}$
@@
(5.13) $\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\sqrt{\frac{\pi}{2}}\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m}{2}\right)\quad\text{for each }a\geq 1/2.\end{split}$
@@
for all $\rho\in(0,\tilde{\delta}_{s})$. Moreover, we may assume that $\tilde{\delta}_{s}$ has been chosen continuously on $s$ since $\{K_{\nu}\}_{\nu\in[3/2,(3+p)/2]}$ is equicontinuous. Let $\tilde{\delta}=\delta\land\min_{s\in[0,1]}\tilde{\delta}_{s}$, then $\delta=\delta(\varepsilon,p,R,a)>0$ and (5.14) holds for all $\rho\in(0,\tilde{\delta})$.
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}H_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}\rho^{p}\sinh^{2m+1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq\lim_{s\nearrow 1}(1-s)(1+\varepsilon)^{2m}\int_{0}^{\frac{\delta}{2a}}\rho^{p+2m}(-\partial_{\rho})\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}H_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=(1+\varepsilon)^{2m}(p+2m)\\ &\quad\times\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}\rho^{p+2m-1}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq\frac{(1+\varepsilon)^{2m}}{(1-\varepsilon)^{2m-1}}(p+2m)\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}H_{s,p,m,a}(\rho)\,\mathrm{d}\rho\\ &=\frac{(1+\varepsilon)^{2m}}{(1-\varepsilon)^{2m-1}}\sqrt{\frac{\pi}{2}}\frac{2^{m}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+2}{2}\right).\end{split}$
@@
$\lim_{s\nearrow 1}\int_{0}^{R}\rho^{p+\beta}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0\quad\text{for each }a\geq 1$
@@
$\begin{split}\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p+\beta}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho=0\end{split}$
@@ Appendix A Auxiliary result
[…]

### v3

@@ 5. Pointwise convergence
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}G_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &\leq(1+\varepsilon)^{2m+1}(p+2m+1)\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p+2m}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}(p+2m+1)\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &=\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}(p+2m+1)\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\\ &=\frac{(1+\varepsilon)^{2m+1}}{(1-\varepsilon)^{2m}}\frac{2^{m}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+3}{2}\right).\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\sqrt{\frac{\pi}{2}}\frac{2^{m-1}}{p}\left(\frac{2}{m-1/2}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m}{2}\right).\end{split}$
@@
(5.13) $\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\sqrt{\frac{\pi}{2}}\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m}{2}\right)\quad\text{for each }a\geq 1/2.\end{split}$
@@
[…]
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}H_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}\rho^{p}\sinh^{2m+1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq\lim_{s\nearrow 1}(1-s)(1+\varepsilon)^{2m}\int_{0}^{\frac{\delta}{2a}}\rho^{p+2m}(-\partial_{\rho})\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}H_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=(1+\varepsilon)^{2m}(p+2m)\\ &\quad\times\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}\rho^{p+2m-1}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq\frac{(1+\varepsilon)^{2m}}{(1-\varepsilon)^{2m-1}}(p+2m)\lim_{s\nearrow 1}(1-s)\int_{0}^{\frac{\delta}{2a}}H_{s,p,m,a}(\rho)\,\mathrm{d}\rho\\ &=\frac{(1+\varepsilon)^{2m}}{(1-\varepsilon)^{2m-1}}\sqrt{\frac{\pi}{2}}\frac{2^{m}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+2}{2}\right).\end{split}$
@@
$\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p+\beta}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0\quad\text{for each }a\geq 1$
@@
$\begin{split}\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p+\beta}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho=0\end{split}$
@@ Appendix A Auxiliary result
Let $p>1$, $r>0$, $u\in C^{2}_{b}(\mathbb{H}^{n})$, and $x\in\mathbb{H}^{n}$. If $p\in(1,\frac{2}{2-s}]$, assume $\nabla u(x)\neq 0$ additionally. If $K:\mathbb{H}^{n}\to\mathbb{R}$ is rotationally symmetric with respect to $x$, that is, $K(\xi)=K(d(x,\xi))$ for all $\xi\in\mathbb{H}^{n}$, and $\int_{\mathbb{H}^{n}}K(\xi)|\xi|^{\alpha}\,d\xi<\infty$, then